OpenStudy (anonymous):

Find the inverse of h(x)=2x+5 h^-1(x)=

OpenStudy (anonymous):

y= (x-5)/2

OpenStudy (anonymous):

y= h(x) x=h^-1y y=2x+5 x=y-5/2 h^-1(y)= y-5/2 or h^-1(x)=x-5/2